-2(x+5)-(3x-8)=3(x-1)
Answer by stanbon(75887)

-2(x+5)-(3x-8)=3(x-1)
-2x-10-3x+8 = 3x-3
-5x-2 = 3x-3
8x = 1
x = 1/8
===============
Cheers,
Stan H.